When the 48 aliens first arrive on Earth they are all wearing either green or blue or both. 27 are wearing blue. The number that is wearing green is 8 more than that.
How many aliens are wearing just green?
So far I got:
8 green
27 both
13 blue
Is this right? My CMT is next week! Please help!

Answer by Edwin McCravy(20064)

No, that's incorrect.  Here's how to do it:



Draw two circles overlapping. One is labeled G for green, and
the other labeled B for blue. That makes 3 regions.
There are x in the region on the left, there are
y in the region in the middle, and z in the region on
the right.

x are wearing just green, y are wearing both green and 
blue, and z are wearing just blue.

Since there are 48 aliens, that tells us that

x + y + z = 48

We are told 

27 are wearing blue.
Notice that some of those 27 are wearing blue only
(these are the z in the region on the far right, and 
the rest of those 27 are in the middle region who also 
are wearing green, so

Since 27 are wearing blue, that tells us that

y + z = 27

Then we are told:

The number that are wearing green is 8 more than that.
8 more than 27 is 27+8 or 35. Notice that some of those 
35 are wearing green only (these are the x in the region 
on the far left, and the rest of those 35 are in the middle 
region who also wear blue, so

x + y = 35

That gives us the system of equations:

(1)    x + y + z = 48
(2)        y + z = 27
(3)    x + y     = 35

Subtract equation (2) from equation (1) term by term
and get

       x         = 21

Subtract equation (3) from equation (1) term by term
and get

               z = 13

Substitute that in equation (2)

           y + z = 27
          y + 13 = 27
               y = 14

So x = 21, y = 14, z = 13.



The question asks how many are just wearing green.
That's the value of x, which is 21.

Edwin